Adjusting Entries

Prepare the necessary adjusting entries at year end of December 31, 2014 for the JS

Corporation for each of the following. No adjusting entries were made during the year.

If no adjustment is needed, state that fact. Round to the nearest dollar.

1. On December 20, 2014, JS received a $4,000 payment from a customer for

services to be rendered early in 2015. Service revenue was credited.

2. On December 1, 2014, JS paid a local radio station $3,000 for 60 radio ads that

were to be aired 20 per month throughout December, January, & February.

Advertising expense was debited.

3. Employee salaries for the month of December totaling $16,000 will be paid on

January 8, 2015.

4. A three year fire insurance policy was purchased on July 1, 2014 for $12,000.

Insurance expense was debited.

5. Depreciation on equipment totaled $15,000 for the year.

6. On August 31, 2014, the company borrowed $60,000 from the bank. A note was

signed with principal and 8% interest to be paid on August 31, 2015.

7. JS began the year with $2,000 in its supplies account. During the year, $6,600 in

supplies were purchased and debited to Supplies Expense. At year end,

supplies costing $3,100 remain on hand.

Adjusting Entries

Prepare the necessary adjusting entries at year end of December 31, 2014 for the JS

Corporation for each of the following. No adjusting entries were made during the year.

If no adjustment is needed, state that fact. Round to the nearest dollar.

1. On December 20, 2014, JS received a $4,000 payment from a customer for

services to be rendered early in 2015. Service revenue was credited.

2. On December 1, 2014, JS paid a local radio station $3,000 for 60 radio ads that

were to be aired 20 per month throughout December, January, & February.

Advertising expense was debited.

3. Employee salaries for the month of December totaling $16,000 will be paid on

January 8, 2015.

4. A three year fire insurance policy was purchased on July 1, 2014 for $12,000.

Insurance expense was debited.

5. Depreciation on equipment totaled $15,000 for the year.

6. On August 31, 2014, the company borrowed $60,000 from the bank. A note was

signed with principal and 8% interest to be paid on August 31, 2015.

7. JS began the year with $2,000 in its supplies account. During the year, $6,600 in

supplies were purchased and debited to Supplies Expense. At year end,

supplies costing $3,100 remain on hand.
Answer #1
JS Corporation
Adjusting Entries
Date Account Debit Credit Calculation
31-Dec Service revenue $        4,000
Unearned Service revenue $        4,000
31-Dec Prepaid Advertising $        2,000 =3,000/3* 2 months
Advertising expense $        2,000
31-Dec Salaries expense $     16,000
Cash $     16,000
31-Dec Prepaid Insurance $     10,000 =12,000/36*(36-6)
Insurance expense $     10,000
31-Dec Depreciation expense $     15,000
Accumulated Depreciation $     15,000
31-Dec Interest expense $        1,600 =60,000*8%*4/12
Interest Payable $        1,600
31-Dec Supplies expense $        5,500 =2,000+6,600-3,100
Supplies $        5,500